internal Rate of Return Method—Two Projects

Munch N’ Crunch Snack Company is considering two possible investments: a delivery truck or a bagging machine. The delivery truck would cost $33,887.7 and could be used to deliver an additional 49,000 bags of pretzels per year. Each bag of pretzels can be sold for a contribution margin of $0.38. The delivery truck operating expenses, excluding depreciation, are $0.52 per mile for 17,000 miles per year. The bagging machine would replace an old bagging machine, and its net investment cost would be $35,662.5. The new machine would require three fewer hours of direct labor per day. Direct labor is $15 per hour. There are 250 operating days in the year. Both the truck and the bagging machine are estimated to have four-year lives. The minimum rate of return is 9%. However, Munch N’ Crunch has funds to invest in only one of the projects.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Compute the internal rate of return for each investment. Use the above table of present value of an annuity of $1. If required, round your present value factor answers to three decimal places and internal rate of return to the nearest percent.

Delivery Truck Bagging Machine
Present value factor        
Internal rate of return     %     %
Answer #1
a
Delivery Truck Bagging Machine
Present value factor 3.465 3.170
Internal rate of return 6% 10%
Workings:
Annual cash flows:
Delivery Truck 9780 =(49000*0.38)-(17000*0.52)
Bagging Machine 11250 =3*15*250
Delivery Truck Bagging Machine
Present value factor =33887.7/9780 =35662.5/11250